10Ω 5Ω sr 2.5mil 20 Ω 12.5 μF
Slope/Displacement Find io(t) if Us (t) 35 sin 4000t V. Suppose that i。(t) = 10 cos(wt + φ), where-360°
Answer #1

on 4 yooo rad/sec203)(20 = (10-toj) 20-20 5A toA tot1o-10 10+10-1o to 5A 107 3.1301-6 3.4, A 3-13 sm (yooot-63.43) 写= cos (10-9) siMe (-yoooヒ十Co{tooot-153.ys) Sh(yooot-63.ys) ニ TOF (yooot-153-yr) 3.13